OpenStudy (anonymous):

need help plz. will MEDAL and FAN Michelle and Maggie are at baseball practice. Michelle throws a ball into the air, and when it drops to a height of 5ft, she hits the ball. The height of the ball is modeled by the graph below where t = time in seconds and h = height of the ball from the ground. A graph of a parabola is shown. The y-intercepts are at approximately negative 0.1, 0 and 2.1, 0. The vertex is located at the approximate coordinate of 1, 22. Maggie is throwing a ball into the air and catching it. The height of Maggie’s ball is modeled by the function h(t) = –16t2 + 48t + 15.

OpenStudy (anonymous):

Which ball goes higher in the air, the ball that is hit or the ball that is thrown? Use complete sentences, and show all work to explain how you determined the height that each ball reaches.

OpenStudy (anonymous):

@dan815 @Preetha @kiamousekia

OpenStudy (anonymous):

Ok well since i already know thats not you in your picture, but its fine anyways let me see...

OpenStudy (anonymous):

The numbers on the x-axis represent the time in seconds. The numbers in the y-axis represent the height in feet. At a height of 0 ft, that's the ground, so the ground lies along the x-axis. The graph is the path of the ball flying in the air. Michelle hits the ball when it was at 5 ft, that's when the time starts (at 0 seconds). So on her graph, it is represented by the y-intercept, (0, 5). The highest point the ball reached is at the vertex, represented by (1, 21), meaning at 1 second, it is 21 ft in the air. The average rate of change is another word for slope. Michelle's slope = (21 - 5) / (1 - 0) = 16 Maggie's slope = (51 - 15) / (1.5 - 0) = 24 Since Maggie has a higher slope, she has a faster rate of change (her ball went up faster).

OpenStudy (anonymous):

does that mean maggies when higher than michelles

OpenStudy (anonymous):

Medal will be appreciated, and yes.

OpenStudy (anonymous):

thanks you